### v1

# A refinement of the Lorentz local field expression with impact on the Clausius-Mossotti and Lorentz-Lorenz models

Abstract
In the 19th century Mossotti and Clausius formulated an expression that relates electrical permittivity of a dielectric to the product of molecular polarizability and number density. Lorenz and Lorentz extended the use of that expression to also cover the dielectric’s refractive index. These expressions successfully predict the change of the dielectric’s permittivity and refractive index as a function of a wide range of its number density, using the consideration that molecular polarizability should not change significantly. However, the derivation of these equations is based on an approximation of the electric field inside the spherical void that models the environment of a single molecule with its own field removed. A switch to the exact field solution extends the validity of the Clausius-Mossotti and Lorentz-Lorenz equations to the higher values of number density which occur in densified dielectrics. It also provides a significant change in the estimates of molecular polarizability.
1 Introduction
The equation
$\boldsymbol{E}_{\mathrm{local}}$ $=\boldsymbol{E}+\boldsymbol{E}_{1}$
$=\boldsymbol{E}+\frac{1}{3}\cdot\frac{\boldsymbol{P}}{\varepsilon_{0}}$
$=\boldsymbol{E}\cdot\left(1+\frac{\varepsilon_{\mathrm{r}}-1}{3}\right)$
$=\boldsymbol{E}\cdot\frac{\varepsilon_{\mathrm{r}}+2}{3}$ (1)
is used in one of its forms in many publications and textbooks, e.g. [2, 8, 4, 5, Kittel], to model the local, homogeneous field inside a spherical void cavity in a linear, isotropic dielectric medium, when that medium is exposed to average static field $\boldsymbol{E}$ and carries polarization $\boldsymbol{P}$. This local field, also known as the Lorentz field, is used to model the electrical environment seen by an approximately spherical individual molecule, with its own field removed (as the molecule cannot be polarized by its own field). Term $\boldsymbol{E}_{1}$ models the contribution from the surrounding molecules to the void left by the removed molecule. The field of that molecule is $\boldsymbol{E}_{\mathrm{self}}=-\boldsymbol{E}_{1}$, so that $\boldsymbol{E}_{\mathrm{local}}+\boldsymbol{E}_{\mathrm{self}}=\boldsymbol{E}$. The local field solution is then used to derive the Clausius-Mossotti [2, 11] equation which links the product $N\alpha$ of molecular polarizability $\alpha$ and number density $N$ to the medium’s relative permittivity $\varepsilon_{\mathrm{r}}$. It is also used to derive the Lorentz-Lorenz [8, 9] equation that relates $N\alpha$ to refractive index $n$.
This paper traces all steps in establishing these equations and shows that eq. (1) is a rough approximation, the exact solution to the local field being
$\boldsymbol{E}_{\mathrm{local}}$ $=\boldsymbol{E}+\boldsymbol{E}_{1}$
$=\boldsymbol{E}+\frac{1}{2\varepsilon_{\mathrm{r}}+1}\cdot\frac{\boldsymbol{P}}{\varepsilon_{0}}$
$=\boldsymbol{E}\cdot\left(1+\frac{\varepsilon_{\mathrm{r}}-1}{2\varepsilon_{\mathrm{r}}+1}\right)$
$=\boldsymbol{E}\cdot\frac{3\varepsilon_{\mathrm{r}}}{2\varepsilon_{\mathrm{r}}+1},$ (2)
as published in the year 1858 by Maxwell [10] for an analogous situation in a magnetic field, and confirmed by several others [12, 6, 14]. A comparison of the two field solutions in fig. 1 shows that the difference is significant.
Figure 1: Local field factors inside void sphere in dielectric, average surrounding field imposed.
Sections 2 and 3 recall the original Clausius-Mossotti and Lorentz-Lorenz models and relates them to the local field expression they are based on. With the exact field solution refined expressions are derived, and application examples are presented.
Section 4 shows how estimates of molecular polarizability change going from the original to the alternative expressions.
Section 5 provides the background for the field solutions of uniform-field spheres based on a particular distribution of the sphere’s surface charge, for different sets of relative permittivity inside and outside the sphere. It then shows how the application of an external field induces such sphere polarization and provides the combined field solutions, for all sets including generic equations which cover all of them. A verification by finite element modeling is given.
2 The Clausius-Mossotti model
A space around the origin of a coordinate system is filled by a vacuum or by a linear and isotropic dielectric. The space carries an average, vertically oriented static field $\boldsymbol{E}$ induced by a faraway source. Placing into this space a small sphere of different material, be it conductive or dielectric (with a different relative permittivity but also linear and isotropic), will induce a $\cos{\theta}$-distributed surface charge on the sphere surface. Outside the sphere, the extra field $\boldsymbol{E}_{2}$ generated by this surface charge has the characteristic of a pure dipole, placed at the origin. Inside, it produces an extra vertical field component $\boldsymbol{E}_{1}$ that either enhances or attenuates the imposed average field.
In case the material outside the sphere is dielectric and inside is void, the total inside field $\boldsymbol{E}_{\mathrm{local}}=\boldsymbol{E}+\boldsymbol{E}_{1}$ is known as the local field or Lorentz field. It is used to model the field experienced by an approximately round molecule as presented to it by the average imposed field $\boldsymbol{E}$ plus the field $\boldsymbol{E}_{1}$ from its neighbors that are close enough to have a significant influence on the field inside the void sphere.
The polarization moment $\boldsymbol{p}$ of the individual molecule is scaled by polarizability $\alpha$:
$\boldsymbol{p}=\alpha\varepsilon_{0}\boldsymbol{E}_{\mathrm{local}},$ (3)
and polarization density (also called electric polarization or just polarization) in a medium with number density $N$ is then
$\boldsymbol{P}=N\alpha\varepsilon_{0}\boldsymbol{E}_{\mathrm{local}}.$ (4)
By definition
$\boldsymbol{P}=(\varepsilon_{\mathrm{r}}-1)\varepsilon_{0}\boldsymbol{E}.$ (5)
If the material is sparse, such as in a gas, the influence of polarized neighbor molecules can be neglected and the local field equated to the imposed average field $\boldsymbol{E}$. Then eq. (4) and (5) combine to the Newton-Drude expression:
$\varepsilon_{\mathrm{r}}=1+N\alpha.$ (6)
Combining eq. (4) and (5) with the approximative solution for $\boldsymbol{E}_{\mathrm{local}}$ (1) gives the original Clausius-Mossotti equation[2]:
$N\alpha=$ $3\cdot\frac{\varepsilon_{\mathrm{r}}-1}{\varepsilon_{\mathrm{r}}+2}$
$\Leftrightarrow\varepsilon_{\mathrm{r}}=$ $\frac{1+2N\alpha/3}{1-N\alpha/3}.$ (7)
Combining eq. (4) and (5) with the exact solution for $\boldsymbol{E}_{\mathrm{local}}$ (2) gives:
$N\alpha=$ $\frac{(\varepsilon_{\mathrm{r}}-1)(2\varepsilon_{\mathrm{r}}+1)}{3\varepsilon_{\mathrm{r}}}$
$\Leftrightarrow\varepsilon_{\mathrm{r}}=$ $\frac{1+3N\alpha+\sqrt{(1+3N\alpha)^{2}+8}}{4}.$ (8)
Figure 2 compares the three models. Experimental data for gases and liquids, taken from [4], shows that the Newton-Drude model is the worst fit. This dataset cannot clearly show if the Clausius-Mossotti (7) model or the one based on the exact local field solution (8) is better. The Clausius-Mossotti model does exhibit an asymptote, with transition to negative values of the relative permittivity, at $N\alpha=3$. This asymptote is known in literature as ”Curie point” [12] or ”Mossotti catastrophe” [3] but has not been supported by experiments, and disappears in the model based on the exact field equation.
Figure 2: Comparing experimental data with three models for relative permittivity versus the product of number density and molecular polarizability.
3 The Lorentz-Lorenz model
With Maxwell’s identity for the refractive index $n=\sqrt{\varepsilon_{\mathrm{r}}\mu_{r}}$, where $\varepsilon_{\mathrm{r}}$ is the relative permittivity for electronic polarization that is applicable to the high frequencies of light, and magnetic permeability $\mu_{r}=1$ for most dielectrics, the Clausius-Mossotti eq. (7) can be written as the Lorentz-Lorenz equation:
$n^{2}=$ $\frac{1+2N\alpha/3}{1-N\alpha/3}$
$\Leftrightarrow N\alpha=$ $3\cdot\frac{n^{2}-1}{n^{2}+2}$ (9)
Application of the exact local field solution changes this to:
$n^{2}=$ $\frac{1+3N\alpha+\sqrt{(1+3N\alpha)^{2}+8}}{4}$
$\Leftrightarrow N\alpha=$ $\frac{(n^{2}-1)(2n^{2}+1)}{3n^{2}}.$ (10)
To compare models (9) and (10), data is taken from Arndt-Hummel [1], reporting measured refractive index of glass compounds with a $\mathrm{TiO_{2}}$ fraction of either $0.9$, $4.7$ or $7.5~{}\mathrm{mol\%}$, or $\mathrm{Na_{2}O}$ fraction of either $20$ or $33~{}\mathrm{mol\%}$, at different degrees of densification. With molecular mass $M_{\mathrm{Si}}=28.085$, $M_{\mathrm{Ti}}=47.867$, and $M_{\mathrm{O}}=15.999~{}\mathrm{g/mol}$, each compound $i$ with $\mathrm{TiO_{2}}$ molar fraction $k_{i}$, molecular mass
$M_{i}=(1-k_{i})\cdot(M_{\mathrm{Si}}+2M_{\mathrm{O}})+k_{i}\cdot(M_{\mathrm{Ti}}+2M_{\mathrm{O}})$ (11)
and density $\rho_{j}$ will have number density:
$N_{i,j}=\frac{N_{A}\cdot\rho_{j}}{M_{i}}$ (12)
where $N_{A}$ is Avogadro’s number. With linear polarization the weighted compound molecular polarizability is:
$\alpha_{\mathrm{c}_{i}}$ $=(1-k_{i})\cdot\alpha_{\mathrm{SiO_{2}}}+k_{i}\cdot\alpha_{\mathrm{TiO_{2}}}$
$\Leftrightarrow\boldsymbol{\alpha}_{\mathrm{c}}$ $=\boldsymbol{K}\boldsymbol{\alpha_{\mathrm{bm}}},$ (13)
where $\boldsymbol{\alpha_{\mathrm{bm}}}$ is a column vector with the polarizabilities of the base materials and $\boldsymbol{K}$ is a matrix with columns $(1-k_{i})$ and $k_{i}$. Inversely, $\boldsymbol{\alpha}_{\mathrm{bm}}$ can be retrieved from a large number of data points with least square fitting using
$\boldsymbol{\alpha}_{\mathrm{bm}}=(\boldsymbol{K}^{T}\boldsymbol{K})^{-1}\boldsymbol{K}^{T}\boldsymbol{\alpha}_{\mathrm{c}}.$ (14)
A similar apporach is used for the compounds containing $\mathrm{Na_{2}O}$ with $M_{\mathrm{Na}}=22.990$
To produce the plot of fig. 3, the data is pre-processed as follows: 1) Eq. (12) is used to calculate the number density from the density numbers in the data. 2) Eq. (9) and (10) are used to convert the measured refraction index data into values of $N\alpha$, respectively for the original Lorentz-Lorenz (LL) model and for the exact local field based (ELF) model. 3) The $N\alpha$ numbers are divided by the number density values found in the first step to give $\boldsymbol{\alpha}_{\mathrm{c}}$ for the different compounds. 4) Eq. 14 fits $\boldsymbol{\alpha}_{\mathrm{bm}}$ to the values of the previous step.
Figure 3: Compound and base material molecular polarizability $\alpha$ translated from refractive index values reported by Arndt-Hummel.
Fig. 3 shows that the data is affected by some measurement noise. Nevertheless, it is clear that the $\alpha_{\mathrm{c}}$ numbers recovered using the Lorentz-Lorenz model (9) show an unexpected tilt where the exact local field based model does not or much less so. This tilt results in higher values of standard deviation between the data points and the expected horizontal lines of constant $\alpha$ for each compound. The absence of tilt for the values recovered with the ELF model already indicates that the investment in a better field model does pay off.
Figure 4: Refractive index versus density with highly densified $\mathrm{SiO_{2}}$ measured by Lobanov[7] and others[13].
Fig. 4 confirms the validity of the refined ELF model. Up to $\rho=3~{}\mathrm{g/cm^{3}}$, data taken from three different sources quickly diverges from the LL1 model (red line, eq, (9)) but tracks the ELF model (green line, eq.(10)) with great accuracy. The LL1 and ELF models are scaled by such $\alpha_{\mathrm{SiO_{2}}}$ that the curves pass through the reference data point of $n_{\mathrm{pristine}}=1.459$ (yellow diamond). Assuming the ELF model is correct with $\alpha_{SiO_{2}}=41.9~{}\mathrm{\AA^{3}}$ and applying that value gives the LL2 model (red dashed line). Error percentages (see legend) underline the better match of the test data to the ELF model.
[…]
4 Absolute values of molecular polarizability
In the previous section, the curve fitting involved using the Lorentz-Lorenz (LL) and exact local field (ELF) based models to recover molecular polarizability $\alpha$ of the base materials [^2] Note: historic publications from before the adoption of the SI system of units - but sometimes even today - often use Gaussian units, where many numbers and equations are off by a factor $4\pi$. In particular, a molecular polarizability of $1~{}\mathrm{\AA^{3}}$ in Gaussian units equals $4\pi~{}\mathrm{\AA^{3}}$ in the SI system used throughout this paper., presented in table 1.
Table 1: Molecular polarizability in Å$\mathrm{\AA}^{3}$ (SI), derived from refractive index.
Masuno Arndt-Hummel Jeziorkowski
$\alpha_{\mathrm{SiO_{2}}}$ $\alpha_{\mathrm{SiO_{2}}}$ $\alpha_{\mathrm{TiO_{2}}}$ $\alpha_{\mathrm{SiO_{2}}}$ $\alpha_{\mathrm{Na_{2}O}}$
LL-model: $37.0$ $36.5$ $75.1$ $36.1$ $32.2$
ELF-model: $41.9$ $42.0$ $102.5$ $43.5$ $34.4$
Clearly the models result in different values for polarizability. As mentioned in section 2, the Clausius-Mossotti and Lorentz-Lorenz models are based on the concept of a molecule being polarized by the local field that reigns inside the otherwise void sphere that represents its volume. The exact local field solution (eq.(2)) is shown in fig. 7(b). Both the field inside the sphere and the one outside equal the undisturbed field shown in fig. 8(b), minus the field in fig. 5(b), with source term $\sigma_{\mathrm{f_{0}}}$ set to $P$, the polarization in the undisturbed field situation. Surface charge $P\cos{\theta}$ induces bound charge of opposite polarity in the dielectric outside the sphere, leaving a reduced net charge. However, eq. (1), used to calculated the Clausius-Mossotti and Lorentz-Lorenz equations, subtracts the field of fig. 5(a) instead, thus omitting the induction of bound charge in the dielectric surface around the sphere.
As a result, even though the Clausius-Mossotti and Lorentz-Lorenz equations have been used to extrapolate relative permittivity and refraction index with reasonable accuracy (even going from gas phase to liquid phase), this extrapolation diverges from actual values when densified materials are considered, as shown in fig. 4 (difference between red line and blue data points). Also, these models underestimate the absolute values of molecular polarizability significantly as shown in table 1 and the difference between the dashed red line and the blue data points in the same graph.
5 Field of uniform-field spheres
Pure dipole
The field of a pure dipole $\boldsymbol{p}$, located at the origin of a spherical coordinate system and aligned with its vertical axis, is defined by voltage
$V=\frac{p}{4\pi\varepsilon_{0}}\cdot\frac{\cos{\theta}}{r^{2}}.$ (15)
Spheroid dipole
A dipole that is formed by distributing charge over the surface of a sphere with a $\sigma_{0}\cos{\theta}$ surface charge density produces that same voltage $V_{2}=V$ outside the sphere if
$p=\sigma_{0}\cdot\frac{4}{3}\pi R^{3},$ (16)
so
$V_{2}=\frac{\sigma_{0}}{\varepsilon_{0}}\cdot\frac{1}{3}\cdot\frac{R^{3}\cos{\theta}}{r^{2}}.$ (17)
Thus, outside this spheroid dipole, the field is as if a pure dipole were placed at its center. This fulfills the boundary condition that far away the spheroid dipole should have a field like a pure dipole. Given the absence of other charge in the environment outside the sphere, this likeness must persist down to the sphere surface.
With $z=r\cos{\theta}$, in a compact mix of Cartesian and spherical coordinates:
$V_{2}=\frac{\sigma_{0}}{\varepsilon_{0}}\cdot\frac{1}{3}\cdot\frac{R^{3}}{r^{3}}\cdot z,$ (18)
Inside the sphere, a homogeneous, vertical field with voltage
$V_{1}=\frac{\sigma_{0}}{\varepsilon_{0}}\cdot\frac{1}{3}\cdot z$ (19)
satisfies the condition that $V_{1}=V_{2}$ at the $r=R$ boundary.
It is important to consider the type of charge that the source terms $p$ and $\sigma_{0}\cos{\theta}$ represent. In eq. (15) to (19), that is the net charge, which may consist of bound charge and free charge. Taking free charge $\sigma_{\mathrm{f}}=\sigma_{\mathrm{f_{0}}}\cos{\theta}$ as source term, the relations become
$V_{2}$ $=\frac{\sigma_{\mathrm{f_{0}}}}{\varepsilon_{0}}\cdot\frac{1}{2\varepsilon_{2}+\varepsilon_{1}}\cdot\frac{R^{3}}{r^{3}}\cdot z,$
$V_{1}$ $=\frac{\sigma_{\mathrm{f_{0}}}}{\varepsilon_{0}}\cdot\frac{1}{2\varepsilon_{2}+\varepsilon_{1}}\cdot z.$ (20)
where $\varepsilon_{1}$ and $\varepsilon_{2}$ denote the relative permittivity of the materials respectively inside and outside the sphere (and these materials are isotropic and respond linearly to the fields applied).
With $\boldsymbol{E}=-\nabla V$ the electric fields outside and inside the sphere become:
$\boldsymbol{E}_{2}$ $=\frac{\sigma_{\mathrm{f_{0}}}}{\varepsilon_{0}}\cdot\frac{1}{2\varepsilon_{2}+\varepsilon_{1}}\cdot\boldsymbol{k}_{2},$
$\boldsymbol{k}_{2}=\frac{2\cos{\theta}\cdot\hat{\boldsymbol{r}}+\sin{\theta}\cdot\hat{\boldsymbol{\theta}}}{(r/R)^{3}},$
$\boldsymbol{E}_{1}$ $=\frac{\sigma_{\mathrm{f_{0}}}}{\varepsilon_{0}}\cdot\frac{-1}{2\varepsilon_{2}+\varepsilon_{1}}\cdot\hat{\boldsymbol{z}}.$ (21)
Note that homogeneous $\boldsymbol{E}_{1}$ and pure-dipole-like $\boldsymbol{E}_{2}$ both scale similarly if $\varepsilon_{1}$ or $\varepsilon_{2}$ is changed; their relative distribution stays the same.
Graphs
Situations with a void sphere are illustrated in fig. 5 and with a dielectric sphere in fig. 6. Fig. 10(a) covers the generic solution. For these situations $\sigma_{\mathrm{f}}=P$, where $P$ is the polarization of the situation depicted in fig. 8(b) where $\varepsilon_{2}=\varepsilon_{1}=4$. The left-hand part of these and following graphs has white background color for $\varepsilon_{r}=1$ and light green for $\varepsilon_{r}>1$. The right-hand part’s color scale, also used in fig. 11, quantifies field strength normalized to the average reference field strength $E$. Grey field lines contain arrows for their direction; their distance is a measure for inverse field strength. In the left half they are crossed by voltage contour lines. Surface charges are illustrated with an unrealistic non-zero thickness representing their local magnitude, opaque red/blue for positive/negative net charge $\sigma$ and transparent red/blue for $\sigma_{\mathrm{f}}$ and $\sigma_{\mathrm{b}}$. The surface charge responsible for $\boldsymbol{E}$ is shown at the top and bottom of the graphs, quite close to the spheres, but should be thought of so far away that the sphere field remains unaffected by its own mirror charge. The graphs also provide equations for $V$, $\boldsymbol{E}$, $\boldsymbol{P}$, $\boldsymbol{D}$, $\sigma$, $\sigma_{\mathrm{f}}$ and $\sigma_{\mathrm{b}}$ to allow comparing the uniform-field sphere solutions from all viewpoints.
Sphere in otherwise average field
Applying an average electric field $\boldsymbol{E}$ to a medium containing a sphere of different characteristics induces a $\cos{\theta}$-modulated surface charge onto its surface. One could say the sphere reacts by adding a dipole-like field to its environment and by changing the homogeneous field strength inside the sphere.
- conductive sphere
If the sphere is conductive (see fig. 9), free charge $\sigma_{\mathrm{f}}=3\varepsilon_{2}\cdot\varepsilon_{0}E\cos{\theta}$ assembles inside at the sphere surface. If $\varepsilon_{2}>1$, the free charge induces bound charge $\sigma_{\mathrm{b}}=-3(\varepsilon_{2}-1)\varepsilon_{2}\cdot\varepsilon_{0}E\cos{\theta}$ at the dielectric surface just around the sphere, which then counters part of the field from the free charge, leaving net charge $\sigma=3\cdot\varepsilon_{0}E\cos{\theta}$. This generates field $\boldsymbol{E}_{1}$ which exactly cancels the imposed average field $\boldsymbol{E}$ leaving $\boldsymbol{E}+\boldsymbol{E}_{1}=\boldsymbol{0}$ inside the sphere. Outside, the net charge generates dipole field $\boldsymbol{E}_{2}$. The combined
$\boldsymbol{E}+\boldsymbol{E}_{2}=E\cdot\left(\hat{\boldsymbol{z}}+\hat{\boldsymbol{k}}_{2}\right)$ (22)
peaks at $3E$ outside at the poles ($z=\pm R$) and is reduced to zero outside at the equator ($z=0$).
- dielectric or void sphere
If the sphere is void (fig. 7) or dielectric (fig. 8), a $\cos{\theta}$-modulated surface charge will still appear on the sphere surface, but it will now consist of bound charge in the dielectric boundary, with $\sigma=\sigma_{\mathrm{b}}=\frac{-3(\varepsilon_{2}-\varepsilon_{1})}{2\varepsilon_{2}+\varepsilon_{1}}\cdot\varepsilon_{0}E\cos{\theta}$ (see eq. (25)). The field will still be a linear combination of the externally imposed average field plus the reaction field due to the (now bound) surface charge, adding the dipole field outside and enhancing or attenuating the vertical homogeneous field inside the sphere.
To find out how much surface charge the field $\boldsymbol{E}$ induces onto the sphere, Gauss’s law for net charge on the sphere surface gives:
$\varepsilon_{0}(E_{2,r}-E_{1,r})$ $=\sigma=\sigma_{\mathrm{f}}+\sigma_{\mathrm{b}}|_{r=R}$
$\sigma_{\mathrm{f}}=0\Rightarrow\varepsilon_{0}(E_{2,r}-E_{1,r})$ $=\sigma_{\mathrm{b}}|_{r=R}$ (23)
and Gauss’s law for free charge on the sphere surface gives:
$D_{2,r}-D_{1,r}=\sigma_{\mathrm{f}}|_{r=R},~{}~{}\sigma_{\mathrm{f}}$ $=0$
$\Rightarrow\varepsilon_{2}\varepsilon_{0}(E_{r}+E_{2,r})-\varepsilon_{1}\varepsilon_{0}(E_{r}+E_{1,r})$ $=0|_{r=R}$ (24)
and (23) and (24) combine to solve:
$\sigma_{\mathrm{b}}$ $=\varepsilon_{0}E\cos{\theta}\cdot\frac{-3(\varepsilon_{2}-\varepsilon_{1})}{2\varepsilon_{2}+\varepsilon_{1}},$
$\boldsymbol{E}+\boldsymbol{E}_{2}$ $=E\cdot\left(\hat{\boldsymbol{z}}-\frac{\varepsilon_{2}-\varepsilon_{1}}{2\varepsilon_{2}+\varepsilon_{1}}\cdot\boldsymbol{k}_{2}\right),$
$\boldsymbol{E}+\boldsymbol{E}_{1}$ $=\boldsymbol{E}\cdot\left(1+\frac{\varepsilon_{2}-\varepsilon_{1}}{2\varepsilon_{2}+\varepsilon_{1}}\right),$ (25)
where $\boldsymbol{k}_{2}$ is the generic dipole outside field multiplier introduced in eq. 21. The generic solution is shown in fig. 10(b). A finite element model with results in fig. 11 confirms the analytical solution shown in fig. 7(b). The corresponding voltages are
$V+V_{2}$ $=-E\cdot\left(1+\frac{\varepsilon_{2}-\varepsilon_{1}}{2\varepsilon_{2}+\varepsilon_{1}}\cdot\frac{R^{3}}{r^{3}}\right)\cdot z,$
$V+V_{1}$ $=-E\cdot\frac{3\varepsilon_{2}}{2\varepsilon_{2}+\varepsilon_{1}}\cdot z,$ (26)
Replacing coordinate in eq. (26) $z$ by $r\cos{\theta}$, electric field strength $-E$ by magnetic field strength $I$, inverse dielectric constants $1/\varepsilon_{1}$ and $1/\varepsilon_{2}$ by ”resistances” $k$ and $k^{\prime}$ respectively, voltages $V+V_{2}$ and $V+V_{1}$ by potentials $P$ and $p_{1}$ respectively, and sphere radius $R$ by $a$:
$P$ $=\left(Ir+A\frac{a^{3}}{r^{2}}\right)\cos{\theta},A=\frac{k-k^{\prime}}{2k+k^{\prime}}I$
$p_{1}$ $=Br\cos{\theta},B=\frac{3k}{2k+k^{\prime}}I$ (27)
which is Maxwell’s exact formulation [10].
Derivation using Legendre/Laplacian
An alternative method that leads to equal results is to solve Gauss’s law $\varepsilon_{0}\nabla\cdot\boldsymbol{E}=\sigma$. By definition $\boldsymbol{E}=-\nabla V$ so $\varepsilon_{0}\nabla\cdot(-\nabla V)=\sigma$ which is also written as $\nabla^{2}V=-\sigma/\varepsilon_{0}$ or $\Delta V=-\sigma/\varepsilon_{0}$, where $\Delta$ is called the Laplacian. In areas where $\sigma=0$ the field solution must satisfy $\Delta V=0$ and Legendre derived that for fields with axial symmetry polynomials in spherical coordinates with voltage
$V(r,\theta)=\sum_{\ell=0}^{\infty}\left(A_{\ell}r^{\ell}+\frac{B_{\ell}}{r^{\ell+1}}\right)P_{\ell}(\cos{\theta})$ (28)
satisfy $\Delta V=0$ for any set of fixed coefficients $A_{\ell}$ and $B_{\ell}$. Calculating a field solution then reverts to finding solutions which satisfy the voltage given at boundaries as well as the Legendre polynomial(s) evaluated at those boundaries, and then relying on the polynomials for the field between the boundaries, where there is no net charge.
For the field solutions in this paper it is enough to evaluate polynomials $P_{0}(\cos{\theta})=1$ and $P_{1}(\cos{\theta})=\cos{\theta}$; the higher terms describe fields for multipoles beyond the simple dipole needed here. Then,
  • $A_{0}=V_{\mathrm{bias}}$, can be set to any arbitrary value, 00 in this paper,
  • $B_{0}=\frac{q}{4\pi\varepsilon_{0}}$ models point charge $q$ placed at the origin, which can also be set to 00 as there are no point charges involved in this paper,
  • $A_{1}=-E_{z}$ models a homogeneous, vertically oriented field $\boldsymbol{E}=E_{z}\hat{\boldsymbol{z}}$, and
  • $B_{1}=\frac{p}{4\pi\varepsilon_{0}}$ models a dipole, placed at the origin, with dipole moment oriented vertically $\boldsymbol{p}=p\hat{\boldsymbol{z}}$.
Spheroid dipole
For the case of impressing free surface charge $\sigma_{\mathrm{f}}=\sigma_{\mathrm{f_{0}}}\cos{\theta}$ onto a sphere of relative permittivity $\varepsilon_{1}$ in medium with $\varepsilon_{2}$, coefficients $A_{1}$ and $B_{1}$ are used outside the sphere and, similarly $C_{1}$ and $D_{1}$ inside the sphere. $A_{1}=0$ since the spheroid dipole will not create an average field. $D_{1}=0$ since there is no dipole to generate a dipole field inside. To find $B_{1}$ and $C_{1}$ the first boundary condition to satisfy is Gauss’s law in matter on the sphere surface:
$\sigma_{\mathrm{f}}$ $=(D_{2,r}-D_{1,r})|_{r=R}$
$\Rightarrow\sigma_{\mathrm{f}}$ $=(\varepsilon_{2}\varepsilon_{0}E_{2,r}-\varepsilon_{1}\varepsilon_{0}E_{1,r})|_{r=R}$
$\Leftrightarrow\sigma_{\mathrm{f}}$ $=\left.\left(\varepsilon_{2}\varepsilon_{0}\cdot-\frac{\partial}{\partial r}\left(B_{1}\frac{\cos{\theta}}{r^{2}}\right)\right)\right|_{r=R}$
$-\left.\left(\varepsilon_{1}\varepsilon_{0}\cdot-\frac{\partial}{\partial r}\left(C_{1}r\cos{\theta}\right)\right)\right|_{r=R}$
$\Rightarrow\sigma_{\mathrm{f}}$ $=\sigma_{\mathrm{f_{0}}}\cos{\theta}=\varepsilon_{2}\varepsilon_{0}\cdot 2B_{1}\frac{\cos{\theta}}{R^{3}}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\varepsilon_{1}\varepsilon_{0}\cdot C_{1}\cos{\theta}$
$\Leftrightarrow\frac{\sigma_{\mathrm{f_{0}}}}{\varepsilon_{0}}$ $=2\varepsilon_{2}\cdot\frac{B_{1}}{R^{3}}+\varepsilon_{1}\cdot C_{1}.$ (29)
The second boundary condition is that the voltages inside and outside must match at $r=R$:
$B_{1}\frac{\cos{\theta}}{R^{2}}$ $=C_{1}Rcos{\theta}$
$\Leftrightarrow B_{1}$ $=C_{1}R^{3}.$ (30)
Eq. (29) and (30) combine to
$B_{1}$ $=\frac{\sigma_{\mathrm{f_{0}}}}{\varepsilon_{0}}\cdot\frac{R^{3}}{2\varepsilon_{2}+\varepsilon_{1}},$
$C_{1}$ $=\frac{\sigma_{\mathrm{f_{0}}}}{\varepsilon_{0}}\cdot\frac{1}{2\varepsilon_{2}+\varepsilon_{1}}.$ (31)
These coefficients confirm the field of eq. (21) derived before.
Sphere in otherwise average field
For the case of impressing average field $\boldsymbol{E}$ in the medium containing a sphere, and using the same coefficients, a similar approach is used. Now, $A_{1}=-E$ since the average outside field has to match the imposed average field. $D_{1}=0$ since also here is no dipole to generate a dipole field inside. Again, the first boundary condition to solve is Gauss’s law in matter at the sphere surface:
$\sigma_{\mathrm{f}}=(D_{2,r}-D_{1,r})|_{r=R}$ $=0$
$\Rightarrow\left.\left(\varepsilon_{2}\varepsilon_{0}(E_{r}+E_{2,r})\right)\right|_{r=R}-$
$\left.\left(\varepsilon_{1}\varepsilon_{0}(E_{r}+E_{1,r})\right)\right|_{r=R}$ $=0$
$\Leftrightarrow\varepsilon_{2}\varepsilon_{0}\cdot\left(E\cos{\theta}-\left.\frac{\partial}{\partial r}\left(B_{1}\frac{\cos{\theta}}{r^{2}}\right)\right|_{r=R}\right)-$
$\varepsilon_{1}\varepsilon_{0}\cdot\left(E\cos{\theta}-\left.\frac{\partial}{\partial r}\left(C_{1}r\cos{\theta}\right)\right|_{r=R}\right)$ $=0$
$\Leftrightarrow(\varepsilon_{2}-\varepsilon_{1})\cdot E\cos{\theta}+\varepsilon_{2}\cdot 2B_{1}\frac{\cos{\theta}}{R^{3}}+$
$\varepsilon_{1}\cdot C_{1}\cos{\theta}$ $=0$
$\Leftrightarrow 2\varepsilon_{2}\cdot\frac{B_{1}}{R^{3}}+\varepsilon_{1}\cdot C_{1}=-(\varepsilon_{2}-\varepsilon_{1})\cdot E$
Again, the second boundary condition is that the voltages inside and outside must match at $r=R$:
$B_{1}\frac{\cos{\theta}}{R^{2}}-Ez$ $=C_{1}Rcos{\theta}-Ez$
$B_{1}$ $=C_{1}R^{3}.$ (33)
Eq. (LABEL:eq:divD_given_E+E12_Legendre) and (33) combine to
$B_{1}$ $=-E\cdot R^{3}\frac{\varepsilon_{2}-\varepsilon_{1}}{2\varepsilon_{2}+\varepsilon_{1}},$
$C_{1}$ $=-E\cdot\frac{\varepsilon_{2}-\varepsilon_{1}}{2\varepsilon_{2}+\varepsilon_{1}}.$ (34)
These coefficients confirm the field of eq. (25) derived before.
6 Conclusion
Two different solutions are found in literature for the local field inside a void spherical cavity that is situated in a linear isotropic dielectric exposed to an otherwise homogeneous electric field. One is an approximation, the other one is exact. This paper establishes the exact solution and fully documents two methods that lead towards it. The exact solution is applied to two formulas that relate a dielectric’s density to its relative permittivity (Clausius-Mossotti) and refractive index (Lorentz-Lorenz) respectively. As a result these parameters can be predicted more accurately which is relevant for highly densified materials. The unexplained ”Curie point” or ”Mossotti catastrophe” disappears. Another result is that estimations of molecular polarizabity, based on measurements of relative permittivity or refractive index interpreted with the original Clausius-Mossotti or Lorentz-Lorenz formulas, can be corrected with corrections that can exceed 10% for normal (non-densified) liquids and solids and . The paper shows the importance of solid bookkeeping when it comes to analyzing the influence of net, free and bound charges. Future work could include corrections of textbooks, use of the improved formulas and applying the refinement of this paper to the Debye equation which also contains a permanent polarization term for polar molecules, but is based on the original Clausius-Mossotti equation.
(a) in vacuum
(b) in dielectric
Figure 5: Void sphere with $\sigma_{\mathrm{f_{0}}}\cos{\theta}$ surface charge imposed.
(a) in vacuum
(b) in dielectric
Figure 6: Dielectric sphere with $\sigma_{\mathrm{f_{0}}}\cos{\theta}$ surface charge imposed.
(a) in vacuum
(b) in dielectric
Figure 7: Void sphere with surrounding average field $\boldsymbol{E}$ imposed.
(a) in vacuum
(b) in same dielectric (so field and polarization are uniform)
Figure 8: Dielectric sphere with surrounding average field $\boldsymbol{E}$ imposed.
(a) in vacuum
(b) in dielectric
Figure 9: Conductive sphere with surrounding average field $\boldsymbol{E}$ imposed.
(a) Any dielectric sphere in any dielectric with $\sigma_{\mathrm{f_{0}}}\cos{\theta}$ surface charge imposed.
(b) Any dielectric sphere in any dielectric with surrounding average field $\boldsymbol{E}$ imposed.
Figure 10: Generic equations for any dielectric material inside/outside.
Figure 11: Verification with finite element model, same situation as fig. 7(b). The color scale is also used for figures 5 - 10.
Acknowledgement
The author thanks Reinoud Lavrijsen, Bas Vermulst, Kurt Vergeer, Ton Backx, Dick Harberts, Michel de Lange and Jan van Dijk for their encouragements and help on this paper.